OpenStudy (anonymous):

LOL FREE MEDAL FOR ANSWER SO EASY! What is the area of a triangle with a base of 2.4 cm and a height of 0.8 cm? A. 0.96 cm² B. 1.6 cm² C. 1.92 cm² D. 3.2 cm²

OpenStudy (anonymous):

who Wants dibs

OpenStudy (blm1):

A=hbb 2

OpenStudy (instagrammodel):

0.96 cm. If you plug in the stuff correctly.

OpenStudy (anonymous):

?

jabez177 (jabez177):

The formula for the area of a triangle is \[\frac{ 1 }{ 2 } b \times h\]

OpenStudy (blm1):

1/2 base x height

OpenStudy (anonymous):

so hoo won Get it?

OpenStudy (blm1):

1.2 x 0.8 =

jabez177 (jabez177):

So we multiply the Base which is 2.4 X 0.8

OpenStudy (anonymous):

can I give it to all of you?

OpenStudy (blm1):

No you can only pick one :)

jabez177 (jabez177):

What is 2.4 x 0.8, @Need_Homework-Help? :)

OpenStudy (anonymous):

oh ._.

jabez177 (jabez177):

You can use a calculator. :)

OpenStudy (instagrammodel):

For me, it's not about the medal.. but about being able to help someone that needs it :)

OpenStudy (anonymous):

i dont need one

OpenStudy (blm1):

You could do either 2.4 x 0.8 / 2 = ? or 1.2 x 0.8 = ?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

so its 1.920

OpenStudy (juliannegirltomboy):

its 0.96 may i have medal but its ok if you pick someone els

jabez177 (jabez177):

@Need_Homework-Help Do you need us to help you? We are not allowed to give direct answers so please, work with us! :D 2.4 x 0.8 = ?

OpenStudy (blm1):

No it is not that. You may use a calculator.

OpenStudy (anonymous):

1.92

OpenStudy (juliannegirltomboy):

you 2.4x0.8=1.92*2=0.96 answer0.96

OpenStudy (juliannegirltomboy):

if its a triangle you always divide by 2

OpenStudy (blm1):

No @Need_Homework-Help don't forget the decimal points :)

OpenStudy (blm1):

@Juliannegirltomboy not x 2 / 2

OpenStudy (anonymous):

where doI insert decimals two sides to the left right?

OpenStudy (blm1):

Like this 2.4 x 0.8

OpenStudy (juliannegirltomboy):

yea it is divided by 2 i know it bc im in collage duh

OpenStudy (blm1):

DO this simply 2.4 x 0.8 divided by 2

OpenStudy (blm1):

You said * 2

OpenStudy (anonymous):

yes .96!

OpenStudy (juliannegirltomboy):

the answer is 0.96

OpenStudy (blm1):

YES! Good job! :)

OpenStudy (anonymous):

lol i didnt divide the 2

OpenStudy (anonymous):

thx so much blm!

OpenStudy (blm1):

LOL XD

OpenStudy (anonymous):

medal 2 u

OpenStudy (juliannegirltomboy):

i meant for the * to be a division sighn

OpenStudy (anonymous):

:D

OpenStudy (blm1):

Ok, I just didn't want @Need_Homework-Help to get confused :D

OpenStudy (anonymous):

is there a way to change my name?